### v1

@@ 3 Phenomenology
In this section, we study the phenomenology of our model. We are interested in the minimal scenario, in the sense of the least number of free parameters, which could explain 1) neutrino masses and mixing, 2) dark matter relic density, and 3) the muon $g-2$ anomaly.
@@ 3.1 Neutrino masses and mixing
where $y_{i}>0$. Again, we have factored out the 1-1 component of $m_{D}$. Finally, the light neutrino mass matrix are given by
@@
Table 2: The best fit values of neutrino oscillation parameters and their $3\sigma$ ranges as determined from the global neutrino data Esteban et al. (2024). Here $m^{2}_{\rm solar}=m_{2}^{2}-m_{1}^{2}$ and $m_{\rm atm}^{2}=m_{3}^{2}-m_{1}^{2}$ ($m_{2}^{2}-m_{3}^{2}$) for normal ordering (inverted ordering).
@@
$J=\text{Im}\left(U_{\mu 3}^{\phantom{*}}U_{e3}^{*}U_{e2}^{\phantom{*}}U_{\mu 2}^{*}\right).$ (31)
The remaining two Majorana phases, in principle, can be deduced from the invariants Im[$(U_{e2}^{\phantom{*}}U_{e1}^{*})^{2}$] and Im[$(U_{e3}^{\phantom{*}}U_{e1}^{*})^{2}$enkins and Manohar (2008). However, both Majorana phases have never been determined experimentally. The values of neutrino masses, mixing angles, and the Dirac phase, as determined from the global analysis of neutrino data, are collected in Tab. 2.
An interesting feature of our set up is that the light neutrino mass matrix is rank 2. As a result, the lightest neutrino mass is vanishing. Hence, the other two non vanishing neutrino masses depend only on $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. In particular, the sum of neutrino masses is given by $\sum_{i}m_{i}=\sqrt{\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ and $\sqrt{\Delta m^{2}_{\rm atm}-\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ for the NO and IO scenarios respectively. The sum of neutrino masses is constrained by cosmological measurement. Combining the Planck 2018 cosmic microwave background data and the baryonic acoustic oscillation measurement result in an upper bound $\sum_{i}m_{i}\leq 130$ meV at 95% confidence level (CL) Aghanim et al. (2020). In addition to the neutrino masses sum, there are two effective masses relevant for laboratory based neutrino experiments. First, the effective mass $m_{\nu_{e}}^{\rm eff}=\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ determines the endpoint of beta decay spectrum. The latest measurement published by KATRIN has put an upper bound of $m_{\nu_{e}}^{\rm eff}\leq 450$ meV at 90% CL Aker et al. (2024). Second the rate for neutrinoless double beta ($0\nu\beta\beta$) decay depends on the effective mass $m_{ee}=|\sum_{i}U_{ei}^{2}m_{i}|$. The strongest upper bound on $m_{ee}$ is provided by the KamLAND-Zen experiment with $m_{ee}\leq 36-156$ meV at 90% CL Abe et al. (2023a).
@@ 3.2 Dark matter phenomonology
In addition to generating neutrino masses, our model also provides a DM candidate which is assumed to be a fermion, $\chi$. In the absence of the neutral scalar mixing, $\theta$, DM interaction with the SM sector is mediated by the ALP, see Eq. (22). Moreover, in the spirit of minimality, the ALP is taken to couple only to muons and photons to explain the muon $g-2$ anomaly, see Sec. 3.3. As a result, assuming $\phi$ and $a$ are heavier than $\chi$, there are only two DM annihilation channels
$\sigma_{\gamma\gamma}v_{\mathrm{rel}}$ $\simeq\frac{1}{4\pi^{3}}\frac{\alpha^{2}C_{\gamma\gamma}^{2}m_{\chi}^{6}}{(m_{a}^{2}-4m_{\chi}^{2})^{2}f_{a}^{4}},$ (32)
$\sigma_{\mu\bar{\mu}}v_{\rm rel}$ $\simeq\frac{8}{\pi}\frac{C_{\mu\mu}^{2}m_{\mu}^{2}m_{\chi}^{4}}{(m_{a}^{2}-4m_{\chi}^{2})^{2}f_{a}^{4}}\sqrt{1-\frac{4m_{\mu}^{2}}{m_{\chi}^{2}}},$ (33)
where $m_{\mu}$ is the mass of the muon. Note that the $\mu\bar{\mu}$ annihilation channel is parametrically suppressed by
$\frac{\sigma_{\mu\bar{\mu}}}{\sigma_{\gamma\gamma}}\sim\frac{\pi^{2}}{\alpha^{2}}\frac{C_{\mu\mu}^{2}}{C_{\gamma\gamma}^{2}}\frac{m_{\mu}^{2}}{m_{\chi}^{2}}.$ (34)
Thus, for $m_{\chi}\gtrsim 1$ TeV, one expects the $\mu\bar{\mu}$ channel to be irrelevant unless there is a large hierarchy in $C_{\mu\mu}$ and $C_{\gamma\gamma}$.
The energy density of DM is precisely measured by the Planck satellite Aghanim et al. (2020) to be
@@
Assuming DM is present in the thermal bath of the early Universe, its present energy density is related to its annihilation cross section. The value of DM energy density can be determined by solving Boltzman equation numerically. However, here we will adopt a simple freeze out approximation, which yields the thermal relic cross section $\langle\sigma v_{\rm rel}\rangle=2.6\times 10^{-26}$ cm3/s. In our numerical analysis, we take ±$\pm 0.4\times 10^{-26}$ cm3/s as the nominal error associate with the approximation. The thermal relic cross section, in turns, constrains the possible range of the ALP couplings $C_{\gamma\gamma}$ and $C_{\mu\bar{\mu}}$. In our numerical analysis, we also verify that the our results based on the freeze out approximation are consistent with micrOMEGAs Belanger et al. (2002); Alguero et al. (2024). To this end, we have implemented the model files in LanHEP Semenov (2016), and then feed them to CalCHEP Belyaev et al. (2013) and micrOMEGAs for numerical computation.
The DM annihilation cross sections in Eqs. (32) and (33) also give rise to detectable signals in gamma ray telescopes. The annihilation channel $\sigma_{\gamma\gamma}v_{\mathrm{rel}}$ leads to a line signal in gamma ray spectrum at energy equals to DM mass. Such a signal has been searched for at the FermiLAT telescope in the low DM mass range, $m_{\chi}\lesssim 1$ TeV Ajello et al. (2016). The heavier DM mass range, $m_{\chi}\leq 100$ TeV, has been searched for by H.E.S.S. Abdallah et al. (2018) and MAGIC Abe et al. (2023b).
3.3 The anomalous magnetic moment of muon
As mentioned before, this model has an ability to explain the muon $g-2$ anomaly. Currently, the combination of BNAL and FNAL run 1-3 results in the discrepancy of Aguillard (2023)
$\Delta a_{\mu}\equiv a_{\mu}^{\rm EXP}-a_{\mu}^{\rm SM}=(249\pm 48)\times 10^{-11},$ (36)
where $a_{\mu}=(g-2)/2$. The discrepancy stands at a $5\sigma$ level and invites new physics interpretation.
@@ 3.3 The anomalous magnetic moment of muon
The new positive contributions to $a_{\mu}$ can be induced by the ALP-photon and ALP-muon couplings
$\mathcal{L}\supset C_{\gamma\gamma}\frac{\alpha}{4\pi}\frac{a}{f_{a}}F_{\mu\nu}\tilde{F}^{\mu\nu}+\frac{C_{\mu\mu}}{f_{a}}(\partial_{\mu}a)\overline{\mu}\gamma^{\mu}\gamma_{5}\mu,$
through the Feynman diagrams shown in Fig. 1. The analytical result for each diagram is given by Buen-Abad et al. (2021); Ganguly et al. (2022)
@@
$\Delta a_{\mu}^{\rm(b)}=$ $-\Bigg{(}\frac{C_{\mu\mu}}{f_{a}}\Bigg{)}\Bigg{(}\frac{C_{\gamma\gamma}}{f_{a}}\Bigg{)}\frac{\alpha}{4\pi^{3}}m_{\mu}^{2}\int_{0}^{1}dx\Bigg{[}(1-x)\Bigg{(}\log{\Big{[}\frac{4\pi f_{a}^{2}}{m_{\mu}^{2}x^{2}+m_{a}^{2}(1-x)}\Big{]}}-\frac{1}{2}\Bigg{)}$
$-\frac{3}{m_{a}^{2}/m_{\mu}^{2}}\Bigg{(}x^{2}\log{\Big{[}\frac{x^{2}+m_{a}^{2}/m_{\mu}^{2}(1-x)}{x^{2}}\Big{]}}\Bigg{)}\Bigg{]},$
@@
It should be noted that diagrams 1 (a) and (c) generate a negative and positive contribution to $\Delta a_{\mu}$ respectively. The contribution of diagram 1 (b) could be either positive or negative depending on the relative sign of $C_{\gamma\gamma}$ and $C_{\mu\mu}$.
@@ 4 Numerical results
For the DM and muon $g-2$ sector, there are 5 free parameters: $m_{\chi}$, $m_{a}$, $f_{a}$, $C_{\gamma\gamma}$ and $C_{\mu\mu}$. Here, we identify the region of parameter space consistent with the observed DM relic abundance which can also explain the muon $g-2$ anomaly. We have also taken into account constraints from gamma ray line searches. We find that in our setup, we need a rather heavy $f_{a}$ scale, $f_{a}\gtrsim 10$ TeV. Fig. 3 shows the allowed parameter space on the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane for the benchmark scenario $m_{a}=m_{\phi}=1.1m_{\chi}$ with $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$, $-1$ [GeV-1] and $f_{a}=10$, 100 TeV respectively. The narrow sky blue region shows the parameter space consistent with DM relic density. The green, teal, and deep blue stripes represent the parameter space which can explain the muon $g-2$ anomaly for $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ [GeV-1] respectively. The dotted gold region is ruled out by the overabundance of DM. The rose, coral, and orange hatched regions are ruled out by gamma ray line searches at FermiLAT Foster et al. (2023), H.E.S.S. Abdallah et al. (2018), and Magic Abe et al. (2023b) respectively. We have also included the projected exclusion region of the upcoming Cherenkov Telescope Array (CTA) Abe et al. (2024) which is shown in the crimson hatched region.
In the case of $f_{a}=10$ TeV, we find that only $C_{\mu\mu}/f_{a}=-1$ [GeV-1] can reproduce the DM density and explain the muon $g-2$ anomaly while remaining consistent with the current gamma ray line searches. However, this scenario would be probed by the upcoming CTA. On the other hand, when $f_{a}=100$ TeV, all $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ can provide the solutions to $g-2$ anomaly and being consistent with DM relic density and gamma ray line constraints. Unfortunately, these scenarios lie beyond the reach of CTA.
Figure 3: The region of parameter space on the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane consistent with the DM relic density (sky blue) and can explain the muon $g-2$ anomaly (green, teal and deep blue for $C_{\mu\mu}/f_{a}=-0.01$, $-0.1$ and $-1$ [GeV-1] respectively). In these plots, $m_{a}=m_{\phi}=1.1m_{\chi}$. The hatched region is ruled out by the gamma ray line search and the dotted region produces an over abundance of DM. In the top (bottom) figure, $f_{a}$ = 10 (100) TeV.
@@ 5 Conclusion and discussion
From our analysis, we find two minimal scenarios for the inverse seesaw mechanism. They both lead to NO neutrinos with vanishing lightest neutrino mass. As a result, the total neutrino mass $\sum_{i}m_{i}$ and the effective masses $m_{\nu_{e}}^{eff}$ and $m_{ee}$ are directly related to $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. For the total neutrino mass, we find $\sum_{i}m_{i}\simeq 58$ meV. This is an order of magnitude below the current Planck constraint. However, it can be probed by the Simon Observatory whose projected sensitivity is $\sum_{i}m_{i}\leq 40$ meV at 95% CL Abitbol et al. (2019). Meanwhile, the effective mass $m_{\nu_{e}}^{eff}$ in our scenarios are around 9 meV, which is an order of magnitude below the projection $m_{\nu_{e}}^{eff}\leq 200$ meV at 90% CL of both the HOLMES and the KATRIN experiments Alpert et al. (2015); Aker et al. (2019). Moreover, we find the effective mass $m_{ee}$ is of order a few meV. This is well below the projected sensitivity of $m_{ee}\leq 13-29$ meV at 90% CL of the LEGEND experiment phase II Abgrall et al. (2017). Thus, the key signature for our neutrino sector is a signal in cosmological measurements ($\sum_{i}m_{i}$) without corresponding signals in laboratory based measurements ($m_{\nu_{e}}^{eff}$ and $m_{ee}$).
In our scenario, the DM density is determined from the annihilation $\chi\bar{\chi}\to\gamma\gamma,\,\mu\bar{\mu}$ through a s-channel ALP exchange. For sufficiently large $m_{\chi}$, the diphoton annihilation channel is expected to dominate over the dimuon channel. As a result, the parameter space is tightly constrained by the gamma ray line searches at gamma ray telescopes. In the benchmark scenario where $m_{a}=m_{\phi}=1.1m_{\chi}$, we find that DM can be a thermal relic for $m_{a}\gtrsim 80$ TeV, see Fig. 3. The ALP couplings that determine DM density also generate a positive shift in the muon $g-2$. We identify the region of parameter space in the $m_{a}-C_{\gamma\gamma}/f_{a}$ plane which could explain the muon $g-2$ anomaly for $C_{\mu\mu}/f_{a}=-0.01,-0.1,-1$ [GeV-1] respectively. Most of these parameter space regions are incompatible with DM being a thermal relic and/or the constraints from gamma ray line searches. In the case where $f_{a}=10$ TeV, only the scenario where $C_{\mu\mu}/f_{a}=-1$ [GeV-1] can satisfy both constraints for $m_{a}\sim 100$ TeV. This scenario can be probed by the upcoming CTA. In the case $f_{a}=100$ TeV, all $C_{\mu\mu}/f_{a}=-0.01,-0.1,-1$ [GeV-1] can satisfy both DM relic density and gamma ray line search constraints. Unlike the $f_{a}=10$ TeV case, these region are out of reach of the CTA.
In our setup, we need $f_{a}\gtrsim 10$ TeV and $m_{a}\gtrsim 100$ TeV to be able to explain both the DM relic density and the muon $g-2$ anomaly. This would put both scalars $a$ and $\phi$ out of reach of the LHC and possibly the planed 100 TeV collider as well. The heavy neutrinos, on the other hand, can be at the scale $f_{a}$. Thus, in principle, they could be probed at the proposed 100 TeV collider such as the Future Circular Collider Bernardi et al. (2022).

### v2

@@ 3 Phenomenology
In this section, we study the phenomenology of our model. We are interested in the minimal scenario, in the sense of the least number of free parameters, which could explain 1) neutrino masses and mixing, 2) dark matter relic density, and 3) the muon $g-2$ measurement consistently.
@@ 3.1 Neutrino masses and mixing
where $y_{i}>0$. Again, we have factored out the 1-1 component of $m_{D}$.
Finally, the light neutrino mass matrix are given by
@@
Table 2: The best fit values of neutrino oscillation parameters and their $3\sigma$ ranges as determined from the global neutrino data Esteban:2024eli . Here $m^{2}_{\rm solar}=m_{2}^{2}-m_{1}^{2}$ and $m_{\rm atm}^{2}=m_{3}^{2}-m_{1}^{2}$ ($m_{2}^{2}-m_{3}^{2}$) for normal ordering (inverted ordering).
@@
$J=\text{Im}\left(U_{\mu 3}^{\phantom{*}}U_{e3}^{*}U_{e2}^{\phantom{*}}U_{\mu 2}^{*}\right)=c_{12}s_{12}c_{23}s_{23}c_{13}^{2}s_{13}\sin{\delta_{\rm CP}}.$ (29)
The remaining two Majorana phases, in principle, can be deduced from the invariants Im[$(U_{e2}^{\phantom{*}}U_{e1}^{*})^{2}$] and Im[$(U_{e3}^{\phantom{*}}U_{e1}^{*})^{2}$enkins:2007ip . However, both Majorana phases have never been determined experimentally. The values of neutrino masses, mixing angles, and the Dirac phase, as determined from the global analysis of neutrino data, are collected in Tab. 2.
An interesting feature of our set up is that the light neutrino mass matrix is rank 2. As a result, the lightest neutrino mass is vanishing. Hence, the other two non vanishing neutrino masses depend only on $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. In particular, the sum of neutrino masses is given by $\sum_{i}m_{i}=\sqrt{\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ and $\sqrt{\Delta m^{2}_{\rm atm}-\Delta m^{2}_{\rm sol}}+\sqrt{\Delta m^{2}_{\rm atm}}$ for the NO and IO scenarios respectively. The sum of neutrino masses is constrained by cosmological measurement. Combining the Planck 2018 cosmic microwave background data and the baryonic acoustic oscillation measurement result in an upper bound $\sum_{i}m_{i}\leq 130$ meV at 95% confidence level (CL) Planck:2018vyg . In addition to the neutrino masses sum, there are two effective masses relevant for laboratory based neutrino experiments. First, the effective mass $m_{\nu_{e}}^{\rm eff}=\sqrt{\sum_{i}|U_{ei}|^{2}m_{i}^{2}}$ determines the endpoint of beta decay spectrum. The latest measurement published by KATRIN has put an upper bound of $m_{\nu_{e}}^{\rm eff}\leq 450$ meV at 90% CL Katrin:2024tvg . Second the rate for neutrinoless double beta ($0\nu\beta\beta$) decay depends on the effective mass $m_{ee}=|\sum_{i}U_{ei}^{2}m_{i}|$. The strongest upper bound on $m_{ee}$ is provided by the KamLAND-Zen experiment with $m_{ee}\leq 36-156$ meV at 90% CL KamLAND-Zen:2022tow .
@@ 3.2 Dark matter phenomonology
In addition to generating neutrino masses, our model also provides a DM candidate which is assumed to be a fermion, $\chi$. In the absence of the neutral scalar mixing, $\theta$, DM interaction with the SM sector is mediated by the ALP, see Eq. (18) and (20). Moreover, this model gives additional contribution apart from the SM one which needs to be constrained by the muon $g-2$ measurement, see Sec. 3.3. As a result, assuming $\phi$ and $a$ are heavier than $\chi$, there are nine DM annihilation channels: one of the nine annihilation processes $\chi\overline{\chi}\to f\bar{f}$. Therefore, the total annihilation cross section reads
$\langle\sigma v_{\rm rel}\rangle=\sum_{f}\langle\sigma_{f\overline{f}}v_{\rm rel}\rangle$ (30)
where
$\langle\sigma_{f\overline{f}}v_{\rm rel}\rangle$ $\simeq\frac{8}{\pi}\frac{C_{ff}^{2}X_{\Phi}^{4}m_{f}^{2}m_{\chi}^{4}}{(m_{a}^{2}-4m_{\chi}^{2})^{2}f_{a}^{4}}\sqrt{1-\frac{m_{f}^{2}}{m_{\chi}^{2}}}.$ (31)
Here $m_{f}$ and $m_{a}$ are the SM fermion and axion masses, respectively. The $C_{ff}$ is the axion-SM fermion coupling and $X_{\Phi}$ is the $f_{a}/v_{\Phi}$ ratio.
The energy density of DM is precisely measured by the Planck satellite Planck:2018vyg to be
@@
Assuming DM is present in the thermal bath of the early Universe, its present energy density is related to its annihilation cross section. The value of DM energy density can be determined by solving Boltzman equation numerically. However, here we will adopt a simple freeze out approximation, which yields the thermal relic cross section $\langle\sigma v_{\rm rel}\rangle=2.6\times 10^{-26}$ cm3/s. In our numerical analysis, we take ±$\pm 0.4\times 10^{-26}$ cm3/s as the nominal error associate with the approximation. The thermal relic cross section, in turns, constrains the possible range of the ALP couplings $C_{ff}$.
3.3 The measurement of muon magnetic moment
The recent report on the measurement of muon magnetic moment Aliberti:2025beg ; Muong-2:2025xyk has been confirmed that the measured value agrees with the SM prediction
$\Delta a_{\mu}\equiv a_{\mu}^{\rm EXP}-a_{\mu}^{\rm SM}=(39\pm 64)\times 10^{-11},$ (33)
where $a_{\mu}=(g-2)/2$. There is no tension between measurement and the SM prediction now. We can use this constraint to find the allowed parameter space on $C_{ff}/f_{a}$.
@@ 3.3 The measurement of muon magnetic moment
In this model, the additional contributions to $a_{\mu}$ can be induced by ALP-SM fermion couplings
$\mathcal{L}\supset\frac{C_{ff}}{f_{a}}(\partial_{\mu}a)\overline{f}\gamma^{\mu}\gamma_{5}f,$
through the Feynman diagrams shown in Fig. 1. The analytical result for each diagram is given by Buen-Abad:2021fwq ; Ganguly:2022imo
@@
$\Delta a_{\mu}^{\rm(b)}=$ $\frac{\alpha}{8\pi^{3}}\left(\frac{C_{ff}}{f_{a}}\right)^{2}m_{\mu}^{2}\Big{(}\big{(}I_{e}+I_{\mu}+I_{\tau}+3I_{d}+3I_{s}+3I_{b}\big{)}$
$-\big{(}3I_{u}+3I_{c}+3I_{t}\big{)}\Big{)},$ (34)
where
@@
The total anomalous magnetic moment of the muon is then given by
$\Delta a_{\mu}=\Delta a_{\mu}^{a}+\Delta a_{\mu}^{b}$ (36)
It should be noted that diagrams 1 (a) generate a negative contribution while diagrams 1 (b) can give either negative or positive ones depending on the axion mass and $C_{ff}$. However, the total muon $a_{\mu}$ changes the sign at the axion mass around 500-600 GeV as shown in Fig.2. In that figure, we made a plot of $\Delta a_{\mu}$ as a function of axion mass for various values of $C_{ff}=\{100,500,1000,5000,10000\}$ to compare the effect of $C_{ff}/f_{a}$ on $\Delta a_{\mu}$. Two horizontal lines correspond to the upper and lower limits recently reported by Aliberti:2025beg ; Muong-2:2025xyk . For $C_{ff}=100$ or $C_{ff}/f_{a}=10^{-2}$, we find that the contribution from new diagram has only a mild effect so that the $\Delta a_{\mu}$ is within the range of measurement within ±$\pm\sigma$ for a range of mass from $\{1,10^{4}\}$ GeV. For the larger $C_{ff}/f_{a}$, the smaller allowed range of mass. In addition, the additional $\Delta a_{\mu}$ changes a sign at around 600 GeV.
Figure 2: The $a_{\mu}$ plot as a function of $m_{a}$ for various values of $C_{ff}$ (see legends). Two horizontal lines correspond to upper and lower limits for recent measurement.
@@ 4 Numerical results
Figure 4: The plot of $C_{ff}/fa$ [GeV]-1 and $m_{\chi}$ [GeV] for $m_{a}=m_{\phi}=1.1m_{\chi}$. The green colour labels the region excluded by latest $g-2$ within ±$\pm\sigma$. The pink colour represents the exclusion limit from the ATLAS measurement of $t\bar{t}$ production. The sky blue region is allowed by thermal relic density. The white and dotted gold represent the under and over abundant regions, respectively. A pair of $f\bar{f}$ label the dominant DM annihilation channel.
For the DM and muon $g-2$ sector, we consider five free parameters: the DM mass ($m_{\chi}$), the axion mass ($m_{a}$), the axion decay constant ($f_{a}$), the scalar mixing parameter ($X_{\Phi}$), and the axion-fermion coupling coefficient ($C_{ff}$). We aim to identify the parameter space consistent with three key observations: 1) the observed DM relic abundance, 2) the measured muon magnetic moment, and 3) the axion-top coupling search conducted by ATLAS.
Fig. 4 illustrates the allowed and excluded regions across various current constraints. The green region is excluded by recent muon $g-2$ measurements. The sky blue region represents the parameter space consistent with the thermal relic density, assuming the freeze-out approximation. The pink region indicates the exclusion limit derived from the ATLAS measurement of $t\bar{t}$ production, which is used to constrain axion-top coupling. This measurement, detailed in Ref. Esser:2023fdo , considered the lepton+jets channel with high-$p_{T}$ top quarks and an integrated luminosity of $\mathcal{L}=139$ fb-1. Regions colored white indicate under-abundant DM, while dotted gold signifies over-abundant regions.
At DM masses less than 1 GeV, DM can annihilate into six primary channels: $e^{+}e^{-}$, $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $\mu^{+}\mu^{-}$, and $c\bar{c}$. Around the tau lepton mass ($m_{\tau}$), annihilation into $\tau^{+}\tau^{-}$ also becomes kinematically allowed. Similarly, as the DM mass increases further and exceeds the bottom quark mass ($m_{b}$) and top quark mass ($m_{t}$), the $b\bar{b}$ and $t\bar{t}$ annihilation channels become kinematically accessible. According to the thermal DM annihilation cross-section, the ratio $C_{ff}/f_{a}$ decreases as the DM mass increases. This trend is evident in Fig. 4; when these heavier quark channels open up, a smaller $C_{ff}/f_{a}$ value is required to satisfy the relic density.
Figures 4(a) and (b) demonstrate that increasing $f_{a}$ necessitates a corresponding increase in $C_{ff}$. This is because the ratio $C_{ff}/f_{a}$ must remain constant for specific values of $m_{a}$, $m_{\phi}$, and $X_{\Phi}$ to maintain the relic density. In these plots, the most significant constraint comes from the ATLAS $t\bar{t}$ search, which excludes DM masses below $180$ GeV for $f_{a}=10^{4}$ GeV and below $4$ TeV for $f_{a}=10^{5}$ GeV. When $X_{\Phi}$ is doubled, as shown in Figures 4(c) and (d), the ratio $C_{ff}/f_{a}$ must be suppressed by a factor of four, given that the annihilation cross-section scales as $\braket{\sigma v}\propto(C_{ff}/f_{a})^{2}X_{\Phi}^{4}$. However, this has only a mild effect on the allowed DM parameter space. After increasing $X_{\Phi}$, the allowed DM mass range slightly expands to include masses up to $150$ GeV for $f_{a}=10^{4}$ GeV and up to $2$ TeV for $f_{a}=10^{5}$ GeV.
@@ 5 Conclusion and discussion
From our analysis, we find two minimal scenarios for the inverse seesaw mechanism. They both lead to NO neutrinos with vanishing lightest neutrino mass. As a result, the total neutrino mass $\sum_{i}m_{i}$ and the effective masses $m_{\nu_{e}}^{eff}$ and $m_{ee}$ are directly related to $\Delta m^{2}_{\rm sol}$ and $\Delta m^{2}_{\rm atm}$. For the total neutrino mass, we find $\sum_{i}m_{i}\simeq 58$ meV. This is an order of magnitude below the current Planck constraint. However, it can be probed by the Simon Observatory whose projected sensitivity is $\sum_{i}m_{i}\leq 40$ meV at 95% CL SimonsObservatory:2019qwx . Meanwhile, the effective mass $m_{\nu_{e}}^{eff}$ in our scenarios are around 9 meV, which is an order of magnitude below the projection $m_{\nu_{e}}^{eff}\leq 200$ meV at 90% CL of both the HOLMES and the KATRIN experiments Alpert:2014lfa ; KATRIN:2019yun . Moreover, we find the effective mass $m_{ee}$ is of order a few meV. This is well below the projected sensitivity of $m_{ee}\leq 13-29$ meV at 90% CL of the LEGEND experiment phase II LEGEND:2017cdu . Thus, the key signature for our neutrino sector is a signal in cosmological measurements ($\sum_{i}m_{i}$) without corresponding signals in laboratory based measurements ($m_{\nu_{e}}^{eff}$ and $m_{ee}$).
In our scenario, the DM density is determined from the annihilation $\chi\bar{\chi}\to f\bar{f}$ through a s-channel ALP exchange. For sufficiently large $m_{\chi}$, the $t\bar{t}$ annihilation channel dominates over other di-fermion channels. In the benchmark scenario where $m_{a}=m_{\phi}=1.1m_{\chi}$, we find that DM can be a thermal relic for $m_{a}\gtrsim 150$ GeV, see Fig. 4. The ALP couplings that determine DM density also generate either a positive or a negative shift in the muon $g-2$. We identify the region of parameter space in the $m_{a}-C_{ff}/f_{a}$ plane which could explain the DM relic density, satisfy the muon $g-2$ and the ATLAS $t\bar{t}$ constraints for $f_{a}/[{\rm GeV}]=10^{4},10^{5}$ and $X_{\Phi}=1,2$, respectively. If neutrino and DM masses are generated through our scenario, the LHC Run 3 and/or HL-LHC data could probe the $m_{\chi}\lesssim 2$ TeV regime for $f_{a}=10^{4}$ GeV scenario.